OpenStudy (anonymous):

g(x) = x5-8x4+28x3-56x2+64x-32 what are the zeros

OpenStudy (anonymous):

Mathematica factored the the problem expression as follows:\[x^5-8 x^4+28 x^3-56 x^2+64 x-32=(x-2)^3 \left(x^2-2 x+4\right) \]\[x=\left\{2,2,2,1+i \sqrt{3},1-i \sqrt{3}\right\} \]The terms containing i are the solutions for x in the equation:\[x^2-2 x+4=0 \]